# Large Language Models in Bioinformatics: A Survey

### v1

@@ 5.2 Protein Design and Engineering
Single-cell sequencing technology, an evolution of transcriptome sequencing, enables the examination of gene expression and transcription levels at the individual cell level Potter (2018). This technology is pivotal for understanding various biological processes at the cellular level, such as disease progression, therapeutic efficacy, and resistance. It can identify relevant cell subpopulations and molecules, holding profound significance for the advancement of bioinformatics. However, the effective and accurate processing and analysis of vast single-cell data sets present a challenge for bioinformaticians. Traditional analysis methods and pipelines are based on tools like Seurat and Scanpy Hao et al. (2024b) Wolf et al. (2018). With the advancement of deep learning, numerous studies have integrated frameworks such as deep learning and large language models with the processing and analysis of single-cell data, achieving significant progress and demonstrating the immense potential of combining these approaches to faster development in the field of single-cell research Molho et al. (2024).
scBERT is a pioneering deep learning model designed for cell type annotation in scRNA-seq data. It adapts the BERT (Bidirectional Encoder Representations from Transformers) framework, originally developed for NLP, to the domain of single-cell transcriptomics. The model leverages a pre-training and fine-tuning paradigm, where it first learns general patterns of gene-gene interactions from large-scale unlabeled scRNA-seq data and then fine-tunes on specific tasks using labeled data to predict cell types Molho et al. (2024).
@@
Foundation models like scBERT, Geneformer, and scGPT have transformed single-cell data analysis through self-supervised pretraining on millions of transcriptomes, enabling robust transfer learning for tasks such as cell annotation, perturbation prediction, and batch correction. Architectural innovations (e.g., performer encoders, rank-based encoding) address computational challenges of high-dimensional data, while task-specific designs improve generalizability and interpretability via attention mechanisms.
@@ 7.1.2 Computational Complexity and Model Efficiency
State-of-the-art LLMs, such as AlphaFold and DNABERT, require massive computational resources for both training and inference. This computational barrier limits accessibility, particularly for research groups with limited infrastructure. Furthermore, relatively longer biological sequences significantly increase memory and processing requirements, making it challenging to apply standard Transformer architectures to genome-scale data Bernard et al. (2025). Efficient methods of model compression and retrieval-augmented need to be further explored to enhance scalability.
@@ 7.2.3 Towards Clinical and Biomedical Applications
Bridging the gap between AI-driven bioinformatics and real-world biomedical applications necessitates further advancements in model validation, regulatory compliance, and ethical considerations. Moving forward, LLM tools require rigorous clinical evaluation and experimental benchmarking to ensure healthcare reliability and safety.Perlis and Fihn (2023).

### v2

@@ 5.2 Protein Design and Engineering
Single-cell sequencing technology, an evolution of transcriptome sequencing, enables the examination of gene expression and transcription levels at the individual cell level Potter (2018). This technology is pivotal for understanding various biological processes at the cellular level, such as disease progression, therapeutic efficacy, and resistance. It can identify relevant cell subpopulations and molecules, holding profound significance for the advancement of bioinformatics. However, the effective and accurate processing and analysis of vast single-cell data sets present a challenge for bioinformaticians. Traditional analysis methods and pipelines are based on tools like Seurat and Scanpy Hao et al. (2024b); Wolf et al. (2018). With the advancement of deep learning, numerous studies have integrated frameworks such as deep learning and large language models with the processing and analysis of single-cell data, achieving significant progress and demonstrating the immense potential of combining these approaches to faster development in the field of single-cell research Molho et al. (2024).
scBERT is a pioneering deep learning model designed for cell type annotation in scRNA-seq data Yang et al. (2022). It adapts the BERT framework, originally developed for NLP, to the domain of single-cell transcriptomics. The model leverages a pre-training and fine-tuning paradigm, where it first learns general patterns of gene-gene interactions from large-scale unlabeled scRNA-seq data and then fine-tunes on specific tasks using labeled data to predict cell types Molho et al. (2024).
@@
Foundation models like scBERT Yang et al. (2022), and scGPT Cui et al. (2024) have transformed single-cell data analysis through self-supervised pretraining on millions of transcriptomes, enabling robust transfer learning for tasks such as cell annotation, perturbation prediction, and batch correction. Architectural innovations (e.g., performer encoders, rank-based encoding) address computational challenges of high-dimensional data, while task-specific designs improve generalizability and interpretability via attention mechanisms.
@@ 7.1.2 Computational Complexity and Model Efficiency
State-of-the-art LLMs, such as AlphaFold and DNABERT, require massive computational resources for both training and inference. This computational barrier limits accessibility, particularly for research groups with limited infrastructure (Table 2 and Table 3). Furthermore, relatively longer biological sequences significantly increase memory and processing requirements, making it challenging to apply standard Transformer architectures to genome-scale data Bernard et al. (2025). Efficient methods of model compression and retrieval-augmented need to be further explored to enhance scalability.
@@ 7.2.3 Towards Clinical and Biomedical Applications
Bridging the gap between AI-driven bioinformatics and real-world biomedical applications necessitates further advancements in model validation, regulatory compliance, and ethical considerations. Moving forward, LLM tools require rigorous clinical evaluation and experimental benchmarking to ensure healthcare reliability and safety Perlis and Fihn (2023).
@@
Model Type Avg. Graphical Memory per Device Avg. Training Duration
Encoder-only ~43 GB ~14 Days
Decoder-only ~46 GB ~5 Days
Encoder–Decoder ~81 GB ~40 Days
Table 2: Statistics of average training duration and graphical memory per device across different LLM architectures. Values are computed based on surveyed models.
@@
Acknowledgements
We want to thank our anonymous AC and reviewers for their feedback. This study was supported by the Chinese University of Hong Kong (CUHK; award numbers 4937025, 4937026, 5501517 and 5501329 to Y.L.) and the IdeaBooster Fund (IDBF23ENG05 and IDBF24ENG06 to Y.L.) and partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region (Hong Kong SAR), China (project no. CUHK 24204023 to Y.L.) and a grant from the Innovation and Technology Commission of the Hong Kong SAR, China (project no. GHP/065/21SZ and ITS/247/23FP to Y.L.). This research was also funded by the Research Matching Grant Scheme at CUHK (award numbers 8601603 and 8601663 to Y.L.) from the Research Grants Council, Hong Kong SAR, China.
Appendix A Appendix
Model Computational Resources Graphical Memory per Device Training Duration
DNABERT 8 * NVIDIA 2080Ti 11 GB
AlphaFold2 Tensor Processing Unit (TPU) v3 32 GiB ~10 Days
RoseTTAFold NVIDIA V100 32 GB ~28 Days
Enformer 64 * TPU v3 32 GiB ~3 Days
ESM-1b
MSA Transformer 32 * NVIDIA V100 32 GB
Fold2Seq 2 * Tesla K80 24 GB
ProGPT2 128 * NVIDIA A100 80 GB ~4 Days
scBERT
ProteinBERT Nvidia Quadro RTX 5000 16 GB 28 Days
ProtTrans NVIDIA V100 16 GB
ProGen TPU v3 32 GiB ~14 Days
ESMFold NVIDIA V100 32 GB
Geneformer 12 * NVIDIA V100 32 GB ~3 Days
HyenaDNA 1 * NVIDIA A100 40 GB ~80 Minutes
Uni-RNA 128 * NVIDIA A100 80 GB
ProGen2 TPU v3 32 GiB
xTrimoPGLM 96 * DGX-A100 320 GB ~150 Days
RNA-MSM 8 * NVIDIA V100 32 GB
TFBert 1 * NVIDIA V100 32 GB ~14 Days
DNAGPT 8 * NVIDIA V100 32 GB ~1 Day
GROVER NVIDIA A100 80 GB
megaDNA NVIDIA A100 + 3090Ti 40 GB (A100) + 24 GB (3090Ti)
Nucleotide Transformer 128 * NVIDIA A100 80 GB ~28 Days
RNAErnie 4 * NVIDIA V100 32 GB ~10 Days
RhoFold+ 8 * NVIDIA A100 80 GB ~30 Days
GPTCelltype
Evo 64 * NVIDIA H100 + 128 * NVIDIA A100 80 GB (H100 + A100)
scFoundation NVIDIA Ampere GPU
scGPT NVIDIA A100 80 GB
AlphaFold3 NVIDIA A100 80 GB
DNABERT2 8 * NVIDIA 2080Ti 11 GB ~14 Days
ESM-DBP 4 * NVIDIA V100 16 GB ~3 Days
RoseTTAFold All-Atom 8 * NVIDIA A6000 48 GB ~8 Days
ProstT5 8 * NVIDIA A100 80 GB ~10 Days
EpiGePT NVIDIA 3090 + NVIDIA 4090 24 GB (3090 + 4090)
RiNALMo 7 * NVIDIA A100 80 GB ~14 Days
ENBED 8 * NVIDIA A100 80 GB ~480 GPU Hours
GPN-MSA 4 * NVIDIA A100 80 GB ~3.5 Hours
ESM-3
Table 3: Quantitative overview of computational costs for selected LLMs applied to bioinformatics. ("" denotes unreported data. 1 GiB = 1.07 GB)